livingswag413:

The ABC Toy Company is creating two similar pieces for a board game, as shown below. Find the value of x that makes the two pieces similar.

livingswag413:

1 attachment
livingswag413:

loll hey guys its me again loll

livingswag413:

i was wondering if one of you guys can lend a hand with this question

justjm:

Use CPCTC to form a proportion and find x \( \frac{DE}{AB} = \frac{EF}{BC}\) \(\frac{12}{6} = \frac{x}{7}\) Find x

livingswag413:

Thank You Jm!

livingswag413:

Thank You Jm!